OpenStudy (anonymous):

Where does the graph of y = 2x2 + 7x + 3 cross the x-axis? (−1 over 2, 0) and (−3, 0) (1 over 2, 0) and (3, 0) (−3 over 2, 0) and (−1, 0) (3 over 2, 0) and (1, 0)

OpenStudy (anonymous):

When y = 0. Solve the following for x:\[2 x^2+7 x+3=0\]\[\left\{x\to -3,x\to -\frac{1}{2}\right\} \]